A motorboat travels 406 km in 7 hours going upstream and 936 km in 9 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?
Answer by TimothyLamb(4379)

---
s = d / t
x = speed of boat in still water
y = speed of the water current
---
up:
x - y = 406/7
x - y = 58
---
down:
x + y = 936/9
x + y = 104
---
system of linear equations:
x - y = 58
x + y = 104
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = speed of boat in still water = 81 kph
y = speed of the water current = 23 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php